OpenStudy (anonymous):

what is the valve of x in the equation 4(2x+1)=27+3(2x-5)? a: 21 b: 9 C: 7.5 D: 4

OpenStudy (anonymous):

Straightforward algebra.\[ 4(2x+1)=27+3(2x-5)\rightarrow 8x+4=27+6x-15\rightarrow 2x=8\rightarrow x=4\]

OpenStudy (anonymous):

4(2x+1) = 27+3(2x-5) 8x+4 = 27+6x-15 8x+4=6x+12 8x+4=6x-6x+12 8x-6x+4=12 2x+4=12 2x+4-4=12 2x=8 2x/2=8/2 X=4 Is there anything that you don’t understand?

OpenStudy (anonymous):

x=4